OpenStudy (decarr432):

If (f + g)(x) = 3x^2 + 2x – 1 and g(x) = 2x – 2, what is f(x)?

OpenStudy (decarr432):

@mathmale

OpenStudy (alivejeremy):

he's not coming

OpenStudy (decarr432):

Fair enough he is probably busy

OpenStudy (decarr432):

Can you help?

OpenStudy (alivejeremy):

Nope im not good at that

OpenStudy (decarr432):

3x2 + 1 3x2 – 1 3x2 + 4 3x2 – 4 x2 – 2

OpenStudy (decarr432):

okay thats anyway

OpenStudy (alivejeremy):

but i'll try

OpenStudy (decarr432):

Oh good :)

OpenStudy (alivejeremy):

nvm ik the answer

OpenStudy (alivejeremy):

a) g^-1(x), the inverse of g(x) Let’s follow a few steps for inverse functions: Put y for the g(x) expression, Switch x for y Solve for y as the inverse function Thus y = 2x + 2 Switching x for y, X = 2y + 2 2y = x – 2 Y = (x-2)/2 So g^-1(x) = (x-2)/2

OpenStudy (alivejeremy):

im 100% sure

OpenStudy (decarr432):

It doesnt match an answer choice lol but good try

OpenStudy (alivejeremy):

b) (fog^-1)(2) means substituting g inverse (g^-1) into f(x) when X = 2 Thus, fog^-1(x); x = 2 fog^-1(x) = 3((x-2)/2)^2) + 1 when x = 2 fog^-1(2) = 3((2-2)/2)^2) + 1 = 0 +1 = 1

OpenStudy (alivejeremy):

is one a answer choice

OpenStudy (alivejeremy):

#1

OpenStudy (decarr432):

I posted the answer choices

OpenStudy (alivejeremy):

3x2 – 1

OpenStudy (decarr432):

Okay thanks

OpenStudy (alivejeremy):

if it's wrong tell me :)))))

OpenStudy (decarr432):

I will